If grass grows at the rate of 6 inches per week, how many feet per hour does it grow?
Answer by checkley77(12844)

.5/24*7=x/1
.5/168=x/1 crosss multiply.
168x=.5
x=.5/168
x=.003 ft. per hour is the growth rate.
Proof
24*7*.003=.5
.5~.5